# V2V-LLM: Vehicle-to-Vehicle Cooperative Autonomous Driving with Multi-Modal Large Language Models

### v1

@@ 8 Additional Planning Results
In the main paper, all experiments use point clouds at a single frame from each CAV as the visual input to the models. In this section, we experiment with feeding visual features from $3$ consecutive frames, the current one and the previous two, as the visual input to the models. Tab. 6 shows the planning results of the new setting together with the original setting from the main paper. In general, using visual inputs from multiple frames improves planning performance. However, the performance gain for V2V-LLM is not as significant as other methods. We conjecture that it is because, even with a single visual input frame, the V2V-LLM model is already capable of inferring future movement directions and max possible movement distances in the next 3 seconds using the point cloud feature maps and the locations of nearby detected objects.
@@
Table 6: V2V-LLM’s planning performance in V2V-QA’s testing split in comparison with baseline methods. L2: L2 distance error. CR: Collision Rate. In each column, the best results are in boldface. and the second-best results are in underline.
9 Evaluation Metric Details
For the grounding questions (Q1, Q2, Q3) and the notable object identification question (Q4), the evaluation metrics are F1 score, precision, and recall. The ground-truth answers and model outputs only contain objects’ center locations. If the center distance between the ground-truth answer and the model output is less than a threshold value, this output is considered as a true positive result. We set the threshold value to be $4$ meters, a typical length of a vehicle.
For the planning question (Q5), the evaluation metrics are L2 errors and collision rates. The ground-truth answers and model outputs only contain $6$ future waypoints, so we calculate the L2 errors on those waypoints. When calculating collision rates, we assume each of the CAVs’ bounding box sizes is $4$ meters in length, $2$ meters in width, and $1.5$ meters in height. We place each CAV’s bounding box at the model’s output future waypoints and calculate the Intersection-over-Union (IOU) between the CAV’s bounding box and every ground-truth object’s annotation bounding box in those future frames. If the IOU is larger than 00, it is considered as a collision.
10 Additional Implementation Details
We use 8 NVIDIA A100-80GB GPUs to train our model. We use the same training hyperparameters from LLaVA-v1.5-7b [25]. We use batch size 32. Adam optimizer is adopted for training with a starting learning rate $52e-$ 5 and a cosine learning rate scheduler with a 3% warm-up ratio. We train our Q1 for one epoch and other QA types for 10 epochs. Our training job can finish within 4 hours.
@@ 11 Additional Dataset Statistics
For the grounding questions (Q1, Q2, Q3) and the notable object identification question (Q4), a QA pair can be categorized into either a positive case or a negative case. If at least one object exists that satisfies the condition specified in the questions, the corresponding QA pair is a positive case. Otherwise, it is a negative case. Tab. 7 summarizes the numbers of QA pairs in each category. This table shows that V2V-QA has sufficient positive and negative data samples in both training and testing splits for each of these QA pairs. The planning task (Q5) is excluded from Tab. 7, as each planning QA pair inherently includes a ground-truth future trajectory in its corresponding answer.
@@
Table 7: Dataset statistics of our V2V-QA’s positive and negative cases in the training and testing splits. Q1: Grounding at a reference location. Q2: Grounding behind a reference object at a location. Q3: Grounding behind a reference object in a direction. Q4: Notable object identification.
We also visualize the distribution of ground truth answer locations relative to the asking CAV for the grounding questions (Q1, Q2, Q3) and the notable object identification question (Q4), as shown in Figs. 7, 8, 9 and 10. In our coordinate system, $x$ is the CAV’s front direction, and $y$ is the CAV’s right direction. For the planning question (Q5), we show the distribution of the ending waypoints in the ground truth answer future trajectories, as shown in Fig. 11. These figures indicate that our V2V-QA  has diverse spatial distributions in the driving scenes.
@@
Figure 7: The distribution of ground-truth answer locations relative to CAV in V2V-QA’s Q1: Grounding at a reference location.
@@
Figure 8: The distribution of ground-truth answer locations relative to CAV in V2V-QA’s Q2: Grounding behind a reference object at a location.
@@
Figure 9: The distribution of ground-truth answer locations relative to CAV in V2V-QA’s Q3: Grounding behind a reference object in a direction.
@@
Figure 10: The distribution of ground-truth answer locations relative to CAV in V2V-QA’s Q4: Notable object identification.
@@
Figure 11: The distribution of ground-truth answer locations relative to CAV in V2V-QA’s Q5: Planning.
@@ 12 Additional Qualitative Results
In this section, we show more qualitative results of our proposed V2V-LLM and other baseline methods in the testing split of V2V-QA’s grounding task in Figs. 12 to 15, notable object identification task in Figs. 16 to 17, and planning task in Figs 18 to 19. The baseline methods include no-fusion, early-fusion, and intermediate-fusion: AttFuse [52], V2X-ViT [51], and CoBEVT [50]. In general, we can observe that our proposed  V2V-LLM’s outputs are closer to the ground-truth answers, in comparison to other baseline methods’ results.
@@
Figure 12: V2V-LLM and baseline methods’ grounding results on V2V-QA’s testing split. Magenta $\circ$: reference locations in questions. Yellow $+$: model output locations. Green $\circ$: ground-truth answers.
@@
Figure 13: V2V-LLM and baseline methods’ grounding results on V2V-QA’s testing split. Magenta $\circ$: reference locations in questions. Yellow $+$: model output locations. Green $\circ$: ground-truth answers.
@@
Figure 14: V2V-LLM and baseline methods’ grounding results on V2V-QA’s testing split. Magenta $\circ$: reference locations in questions. Yellow $+$: model output locations. Green $\circ$: ground-truth answers.
@@
Figure 15: V2V-LLM and baseline methods’ grounding results on V2V-QA’s testing split. Magenta $\circ$: reference locations in questions. Yellow $+$: model output locations. Green $\circ$: ground-truth answers.
@@
Figure 16: V2V-LLM and baseline methods’ notable object identification results on V2V-QA’s testing split. Magenta curve: planned future trajectories in questions. Green $\circ$: ground-truth notable object locations. Yellow $+$: model identification outputs.
@@
Figure 17: V2V-LLM and baseline methods’ notable object identification results on V2V-QA’s testing split. Magenta curve: planned future trajectories in questions. Green $\circ$: ground-truth notable object locations. Cyan $\times$: model identification outputs.
@@
Figure 18: V2V-LLM and baseline methods’ planning results on V2V-QA’s testing split. Green curve: future trajectories in ground-truth answers. Green $\circ$: ending waypoints in ground-truth answers. Yellow curve: model planning outputs. Yellow $\times$: ending waypoints in model outputs.
@@
Figure 19: V2V-LLM and baseline methods’ planning results on V2V-QA’s testing split. Green curve: future trajectories in ground-truth answers. Green $\circ$: ending waypoints in ground-truth answers. Cyan curve: model planning outputs. Cyan $\times$: ending waypoints in model outputs.
@@ 13 Limitation
Fig. 20 shows failure cases of V2V-LLM’s planning results on V2V-QA’s testing split. In a few frames, the model generates future trajectories in the lane of the opposite traffic direction. A potential solution and future work is to include HD map information as additional input to the model. Currently, this approach is not feasible because the base dataset V2V4Real [53] has not released its HD map to the public.
Figure 20: Failure cases of V2V-LLM’s planning results on V2V-QA’s testing split. Green curve: future trajectories in ground-truth answers. Green $\circ$: ending waypoints in ground-truth answers. Yellow curve and Cyan curve: model planning outputs corresponding to CAV_EGO and CAV_1, respectively. Yellow $\times$ and Cyan $\times$: ending waypoints in model outputs corresponding to CAV_EGO and CAV_1, respectively.

### v3

@@ 10 Planning Results with Temporal Inputs
In the main paper, all experiments use point clouds at a single frame from each CAV as the visual input to the models. In this section, we experiment with feeding visual features from $3$ consecutive frames, the current one and the previous two, as the visual input to the models. Tab. 11 shows the planning results of the new setting together with the original setting from the main paper. In general, using visual inputs from multiple frames improves planning performance.
@@
Table 11: V2V-LLM’s planning performance in V2V-QA’s V2V-split in comparison with baseline methods. L2: L2 distance error. CR: Collision rate. In each column, the best results are in boldface. and the second-best results are in underline.
11 Detailed Ablation Results
Tab. 12 shows the detailed ablation results when using only the scene-level features or only the object-level features as input to our V2V-LLM. Both types of input features contribute to the final performance, and object-level features are easier for LLM to digest. Training from scratch achieves worse performance, meaning that pre-training with LLaVA’s VQA tasks improves our V2V-LLM’s performance in V2V-QA.
Method Q1 Q2 Q3 QGr Q4 Q5 Comm (MB) $\downarrow$
Scene-level only 69.9 74.9 65.5 15.4 19.9 12.6 17.9 26.9 13.5 34.4 43.2 40.2 46.7 7.21 15.55 0.4008
Object-level only 69.0 80.9 60.1 26.9 34.7 21.9 17.6 18.3 16.9 37.8 52.6 57.3 48.6 5.24 7.78 0.0068
Scratch 67.6 77.6 60.0 26.5 26.4 26.5 17.2 16.4 18.2 37.1 49.3 52.7 46.3 6.30 5.01 0.4068
V2V-LLM (ours) 70.0 80.1 62.2 30.8 36.3 26.7 21.2 21.5 20.8 40.7 59.7 61.9 57.6 4.99 3.00 0.4068
Table 12: Ablation study in V2V-QA’s V2V-split. Q1: Grounding at a reference location. Q2: Grounding behind a reference object at a location. Q3: Grounding behind a reference object in a direction. QGr: Average of grounding (Q1, Q2, and Q3). Q4: Notable object identification. Q5: Planning. P: Precision. R: Recall. L2: L2 distance error. CR: Collision rate. Comm: Communication cost.
@@ 12 Additional Dataset Statistics
For the grounding questions (Q1, Q2, Q3) and the notable object identification question (Q4), a QA pair can be categorized into either a positive case or a negative case. If at least one object exists that satisfies the condition specified in the questions, the corresponding QA pair is a positive case. Otherwise, it is a negative case. Tabs. 13 and 14 summarizes the numbers of QA pairs in each category, for V2V-QA’s V2V-split and V2X-split respectively. This table shows that V2V-QA has sufficient positive and negative data samples in both training and testing sets for each of these QA pairs. The planning task (Q5) is excluded from Tabs. 13 and 14, as each planning QA pair inherently includes a ground-truth future trajectory in its corresponding answer.
@@
Table 13: Dataset statistics of our V2V-QA’s V2V-split on positive and negative samples.
QA type Train-Pos Train-Neg Test-Pos Test-Neg Total
Q1 247447 247843 62332 66379 624001
Q2 84005 83689 18297 16936 202927
Q3 14346 14394 3421 3044 35205
Q4 4624 1650 1172 536 7982
Total 350422 347576 85222 86895 870115
Table 14: Dataset statistics of our V2V-QA’s V2X-split on positive and negative samples.
We also visualize our V2V-split distribution of ground truth answer locations relative to the asking CAV for the grounding questions (Q1, Q2, Q3) and the notable object identification question (Q4), as shown in Figs. 7, 8, 9 and 10. In our coordinate system, $x$ is the CAV’s front direction, and $y$ is the CAV’s right direction. For the planning question (Q5), we show the distribution of the ending waypoints in the ground truth answer future trajectories, as shown in Fig. 11. We visualize the location distribution of V2V-QA’s V2X-splitin Figs. 12, 13, 14, 15 and 16. These figures indicate that our V2V-QA has diverse spatial distributions in the driving scenes. Compared to NuScenes [3], our  V2V-QA has larger ranges and standard deviations of the ground-truth ending waypoints, as shown in Tab. 15. Therefore, the planning task in our  V2V-QA is more challenging.
Dataset x: forward y: right
min max std min max std
NuScenes -0.9 39.7 10.4 -11.0 11.1 1.9
V2V-QA (ours) -2.1 177.0 28.1 -24.3 12.0 2.4
Table 15: Ranges and standard deviations of ground-truth ending waypoints.
@@
Figure 7: The distribution of ground-truth answer locations relative to CAV in V2V-QA’s V2V-split Q1: Grounding at a reference location.
@@
Figure 8: The distribution of ground-truth answer locations relative to CAV in V2V-QA’s V2V-split Q2: Grounding behind a reference object at a location.
@@
Figure 9: The distribution of ground-truth answer locations relative to CAV in V2V-QA’s V2V-split Q3: Grounding behind a reference object in a direction.
@@
Figure 10: The distribution of ground-truth answer locations relative to CAV in V2V-QA’s V2V-split Q4: Notable object identification.
@@
Figure 11: The distribution of ground-truth answer locations relative to CAV in V2V-QA’s V2V-split Q5: Planning.
(a) x (meters)
(b) y (meters)
(c) distance (meters)
(d) angle (degrees)
Figure 12: The distribution of ground-truth answer locations relative to CAV in V2V-QA’s V2X-split Q1: Grounding at a reference location.
(a) x (meters)
(b) y (meters)
(c) distance (meters)
(d) angle (degrees)
Figure 13: The distribution of ground-truth answer locations relative to CAV in V2V-QA’s V2X-split Q2: Grounding behind a reference object at a location.
(a) x (meters)
(b) y (meters)
(c) distance (meters)
(d) angle (degrees)
Figure 14: The distribution of ground-truth answer locations relative to CAV in V2V-QA’s V2X-split Q3: Grounding behind a reference object in a direction.
(a) x (meters)
(b) y (meters)
(c) distance (meters)
(d) angle (degrees)
Figure 15: The distribution of ground-truth answer locations relative to CAV in V2V-QA’s V2X-split Q4: Notable object identification.
(a) x (meters)
(b) y (meters)
(c) distance (meters)
(d) angle (degrees)
Figure 16: The distribution of ground-truth answer locations relative to CAV in V2V-QA’s V2X-split Q5: Planning.
@@ 13 Additional Qualitative Results
We show more qualitative results of our proposed V2V-LLM and other baseline methods in the testing set of V2V-QA’s grounding task in Figs. 17 to 20, notable object identification task in Figs. 21 to 22, and planning task in Figs 23 to 24. The baseline methods include no-fusion, early-fusion, and intermediate-fusion: AttFuse [52], V2X-ViT [51], and CoBEVT [50]. Results of V2X-split can be seen in Figs. 25 to 31. In general, our proposed  V2V-LLM’s outputs are closer to the ground-truth answers, in comparison to other baseline methods’ results.
@@
Figure 17: V2V-LLM and baseline methods’ grounding results on V2V-QA’s V2V-split testing set. Magenta $\circ$: reference locations in questions. Yellow $+$: model output locations. Green $\circ$: ground-truth answers.
@@
Figure 18: V2V-LLM and baseline methods’ grounding results on V2V-QA’s V2V-split testing set. Magenta $\circ$: reference locations in questions. Yellow $+$: model output locations. Green $\circ$: ground-truth answers.
@@
Figure 19: V2V-LLM and baseline methods’ grounding results on V2V-QA’s V2V-split testing set. Magenta $\circ$: reference locations in questions. Yellow $+$: model output locations. Green $\circ$: ground-truth answers.
@@
Figure 20: V2V-LLM and baseline methods’ grounding results on V2V-QA’s V2V-split testing set. Magenta $\circ$: reference locations in questions. Yellow $+$: model output locations. Green $\circ$: ground-truth answers.
@@
Figure 21: V2V-LLM and baseline methods’ notable object identification results on V2V-QA’s V2V-split testing set. Magenta curve: planned future trajectories in questions. Green $\circ$: ground-truth notable object locations. Yellow $+$: model identification outputs.
@@
Figure 22: V2V-LLM and baseline methods’ notable object identification results on V2V-QA’s V2V-split testing set. Magenta curve: planned future trajectories in questions. Green $\circ$: ground-truth notable object locations. Cyan $\times$: model identification outputs.
@@
Figure 23: V2V-LLM and baseline methods’ planning results on V2V-QA’s V2V-split testing set. Green curve: future trajectories in ground-truth answers. Green $\circ$: ending waypoints in ground-truth answers. Yellow curve: model planning outputs. Yellow $\times$: ending waypoints in model outputs.
@@
Figure 24: V2V-LLM and baseline methods’ planning results on V2V-QA’s V2V-split testing set. Green curve: future trajectories in ground-truth answers. Green $\circ$: ending waypoints in ground-truth answers. Cyan curve: model planning outputs. Cyan $\times$: ending waypoints in model outputs.
Figure 25: V2V-LLM and baseline methods’ grounding results on V2V-QA’s V2X-split testing set. Magenta $\circ$: reference locations in questions. Yellow $\times$: model output locations. Green $\circ$: ground-truth answers.
Figure 26: V2V-LLM and baseline methods’ grounding results on V2V-QA’s V2X-split testing set. Magenta $\circ$: reference locations in questions. Yellow $\times$: model output locations. Green $\circ$: ground-truth answers.
Figure 27: V2V-LLM and baseline methods’ grounding results on V2V-QA’s V2X-split testing set. Magenta $\circ$: reference locations in questions. Yellow $\times$: model output locations. Green $\circ$: ground-truth answers.
Figure 28: V2V-LLM and baseline methods’ notable object identification results on V2V-QA’s V2X-split testing set. Magenta curve: planned future trajectories in questions. Green $\circ$: ground-truth notable object locations. Yellow $\times$: model identification outputs.
Figure 29: V2V-LLM and baseline methods’ notable object identification results on V2V-QA’s V2X-split testing set. Magenta curve: planned future trajectories in questions. Green $\circ$: ground-truth notable object locations. Cyan $\times$: model identification outputs.
Figure 30: V2V-LLM and baseline methods’ planning results on V2V-QA’s V2X-split testing set. Green curve: future trajectories in ground-truth answers. Green $\circ$: ending waypoints in ground-truth answers. Yellow curve: model planning outputs. Yellow $\times$: ending waypoints in model outputs.
Figure 31: V2V-LLM and baseline methods’ planning results on V2V-QA’s V2X-split testing set. Green curve: future trajectories in ground-truth answers. Green $\circ$: ending waypoints in ground-truth answers. Cyan curve: model planning outputs. Cyan $\times$: ending waypoints in model outputs.
@@ 14 Limitation
Fig. 32 shows failure cases of V2V-LLM’s planning results on V2V-QA’s testing set. In a few frames, the model generates future trajectories in the lane of the opposite traffic direction. A potential solution and future work is to include HD map information as additional input to the model. Currently, this approach is not feasible because the base dataset V2V4Real [53] has not released its HD map to the public.
Figure 32: Failure cases of V2V-LLM’s planning results on V2V-QA’s testing set. Green curve: future trajectories in ground-truth answers. Green $\circ$: ending waypoints in ground-truth answers. Yellow curve and Cyan curve: model planning outputs corresponding to CAV_EGO and CAV_1, respectively. Yellow $\times$ and Cyan $\times$: ending waypoints in model outputs corresponding to CAV_EGO and CAV_1, respectively.